(5.2.15) Assume that a procedure yields a binomial distribution with n=5 trials and a probability of success of p=0.50. Use a binomial probability table to find the probability that the number of successes x is exactly 2.
P(2)=
1
Expert's answer
2020-11-24T17:14:30-0500

Solution:

"P(X=2)=_2C_5\times p^2 \times (1-p)^3=10\times 0.5^2 \times 0.5^3=0.3125"

Answer: P(2)=0.3125